# JETSCAPE Collaboration

### v1

@@ V.1 Inclusive jet substructure
In this subsection, we present the results of groomed observables for inclusive jets. The initial hard processes for the simulations were generated using the pythiagun module with HardQCD:all=on.
@@ V.1.1 Relative transverse momentum of jet splittings
Figure 25: Distributions of the relative transverse momentum of jet splittings, $k_{T,g}$ for charged jets with $8060<p^{\mathrm{ch,jet}}_{T}<$ 80 GeV and $|\eta_{\mathrm{ch,jet}}|<0.7$, using $R=0.2$ in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, and the ratios for different grooming algorithms: Soft Drop with $z_{\mathrm{cut}}=0.2$ and $\beta=0$ (left), and dynamical grooming with $a=1.0$ (right). The solid lines and circles with statistical error bars show the results from vacuum matter of jetscape and the experimental data from the ALICE Collaboration [56], respectively. The bands indicate the systematic uncertainties of the experimental data.
Figure 26: Ratios of $k_{T,g}$ distributions between Pb-Pb and $p$-$p$ collisions at $\sqrt{s_{NN}}=5.02$ TeV for charged jets with $8060<p^{\mathrm{ch,jet}}_{T}<$ 80 GeV and $|\eta_{\mathrm{ch,jet}}|<0.7$, using $R=0.2$ for different centralities, $0\%$$10\%$ (left) and $30\%$$50\%$ (right), and grooming algorithms, Soft Drop with $z_{\mathrm{cut}}=0.2$ and $\beta=0$ (upper) and dynamical grooming with $a=1.0$ (lower). The solid lines and squares with statistical error bars show the results from matter+lbt of jetscape and the experimental data from the ALICE Collaboration [56], respectively. The bands indicate the systematic uncertainties of the experimental data.
The distributions of $k_{T,g}$ obtained from Soft Drop and dynamical grooming [64], in $p$-$p$ collisions are shown in Fig. 25, and their modifications in Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV are presented in Fig. 26. The results from our jetscape simulations are compared with the experimental results from ALICE [56].
The Soft Drop and dynamical grooming results from the JETSCAPE PP19 tune show consistency with experimental data for the $k_{T,g}$ distribution in $p$-$p$ collisions, remaining within the uncertainty range. For Pb-Pb collisions, the modification in the distributions aligns with the inclusive full jet results presented in the main text, demonstrating a monotonically decreasing behavior as a function of $k_{T,g}$, which is primarily governed by selection bias by the $p_{T}$ cut and the loss of $p_{T,2}$. They successfully capture the experimental trends, while dynamical grooming results exhibit slight over-suppression. Notably, the Soft Drop results are in excellent agreement with experimental data within the uncertainties.
@@ V.1.2 Groomed Jet Mass
Figure 27: Distributions of the ratio of the Soft Drop groomed jet mass to the jet transverse momentum, $m_{g}/p^{\mathrm{jet}}_{T}$, for full jets with $180160<p^{\mathrm{jet}}_{T}<$ 180 GeV and $|\eta_{\mathrm{jet}}|<1.3$, using $R=0.4$ in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV for different grooming parameters: $z_{\mathrm{cut}}=0.1$ and $\beta=0$ (left), and $z_{\mathrm{cut}}=0.5$ and $\beta=1.5$ (right). The solid lines and circles with statistical error bars show the results from vacuum matter of jetscape and the unfolded experimental data from the CMS Collaboration [45], respectively. The bands indicate the systematic uncertainties of the experimental data.
@@
Figure 28: Distributions of the Soft Drop groomed mass, $m_{g}$ for charged jets with $R=0.2$ and $|\eta_{\mathrm{ch,jet}}|<0.7$ in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, and the ratios for different $p^{\mathrm{ch,jet}}_{T}$ range. The Soft Drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$. The solid lines and circles with statistical error bars show the results from vacuum matter of jetscape and the experimental data from the ALICE Collaboration [57], respectively. The bands indicate the systematic uncertainties of the experimental data.
@@
Figure 29: Ratios of Soft Drop $m_{g}$ distributions between $0\%$$10\%$ Pb-Pb and $p$-$p$ collisions at $\sqrt{s_{NN}}=5.02$ TeV for charged jets with $R=0.2$ and $|\eta_{\mathrm{ch,jet}}|<0.7$ for different $p^{\mathrm{ch,jet}}_{T}$ range. The Soft Drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$. The solid lines and squares with statistical error bars show the results from matter+lbt of jetscape and the experimental data from the ALICE Collaboration [57], respectively. The bands indicate the systematic uncertainties of the experimental data.
@@
Here, we present a comparison of our results for Soft Drop groomed jet mass $m_{g}$ with experimental results. The comparison of the unfolded $m_{g}/p_{T}$ distribution for full jets with $R=0.4$ in $p$-$p$ collisions from CMS [45] is shown in Fig. 27, while the comparison of the $m_{g}$ distribution for charged jets with $R=0.2$ in $p$-$p$ collisions from ALICE [57] is shown in Fig. 28. The results for $p$-$p$ collisions are in agreement with the experimental data within the uncertainty range, except for regions with significantly low jet counts, such as the tails of the distributions. It should be noted that, while our calculations are based on Monte Carlo simulations, the masses of jet constituents were adjusted to align with experimental methodologies: massless for comparison with CMS data and the charged pion mass for comparison with ALICE data. This adjustment, while necessary to match experimental analysis techniques, has a non-negligible impact on the results, leading to noticeable changes in the distributions.
@@
Figure 29 shows the modifications of the Soft Drop groomed jet mass distributions in $0\%$$10\%$ Pb-Pb collisions, compared with the ALICE data [57]. As with the inclusive full jet results presented in the main text, the behavior is dominated by the $p^{\mathrm{jet}}_{T}$ cut-induced selection bias and the loss of prong mass, leading to suppression at large $m_{g}$. The results capture the experimental trends and agree with the data within the uncertainty range across most $m_{g}$ regions. A slight enhancement observed in some $p^{\mathrm{jet}}_{T}$ ranges for the largest $m_{g}$ bin is primarily attributed to the very low jet counts in the tail of the distribution, particularly in $p$-$p$ collisions.
@@ V.2 Jet splitting radius of $\gamma$-tagged jet
Figure 30: Distributions of the Soft Drop groomed mass, $r_{g}$ for $\gamma$-tagged jets with $R=0.2$ in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, and the ratios for $x_{J\gamma}>0.8$ (left) and $x_{J\gamma}>0.4$ (right). Leading full jets within $|\eta_{\mathrm{jet}}|<2$ and $|\phi_{\mathrm{jet}}-\phi_{\gamma}|>2\pi/3$ are paired with isolated photons satisfying $p^{\gamma}_{T}>100~{}\mathrm{GeV}$ and $|\eta_{\gamma}|<1.44$ as the trigger condition. The Soft Drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$. The solid lines and circles with statistical error bars show the results from vacuum matter of jetscape and the experimental data from the CMS Collaboration [55], respectively. The bands indicate the systematic uncertainties of the experimental data.
Figure 31: Ratios of Soft Drop $r_{g}$ distributions between $0\%$$30\%$ Pb-Pb and $p$-$p$ collisions at $\sqrt{s_{NN}}=5.02$ TeV. Full jets of highest $p^{\mathrm{jet}}_{T}$ within $|\eta_{\mathrm{jet}}|<2$ and $|\phi_{\mathrm{jet}}-\phi_{\gamma}|>2\pi/3$ are paired with isolated photons satisfying $p^{\gamma}_{T}>100~{}\mathrm{GeV}$ and $|\eta_{\gamma}|<1.44$ as the trigger condition. The Soft Drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$. The solid lines and squares with statistical error bars show the results from matter+lbt of jetscape and the experimental data from the CMS Collaboration [55], respectively. The bands indicate the systematic uncertainties of the experimental data.
Figures 30 and 31 present comparisons with CMS data [55] for $r_{g}$ distributions in $p$-$p$ collisions and their modifications in $0\%$$30\%$ Pb-Pb collisions, respectively. Results are shown for two lower cuts on $x_{J\gamma}$: $0.8$ and $0.4$. For $p$-$p$ collisions, the distributions obtained from the vacuum matter calculations are narrower than the experimental data for both $x_{J\gamma}$ cut values. This behavior is consistent with the trends observed in other Monte Carlo model calculations compared in Ref. [55].
For the Pb-Pb results, at $x_{J\gamma}>0.8$, suppression at large $r_{g}$ region, attributed to selection bias, is observed, similar to the inclusive jet results presented in the main text and in our previous study [74], as well as the $\gamma$-tagged jet results with large $x_{J\gamma}$ cuts in the main text. While there is a sizable quantitative difference compared to the experimental data, the results qualitatively reproduce the observed trends.
@@ VI Summary
In this work, we investigated medium modifications in jet substructure observables from Soft Drop grooming, including the jet splitting momentum fraction ($z_{g}$), splitting radius ($r_{g}$), relative transverse momentum of splittings ($k_{T,g}$), and groomed jet mass ($m_{g}$), in $0\%$$10\%$ Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV. These analyses utilized the matter$+$lbt multistage jet evolution model within the jetscape framework, using the JETSCAPEv3.5 AA22 tune [89], as employed in previous studies [90, 74]. In particular, this effort should be compared to the closely related earlier companion paper Ref. [93], which studied medium effects on $\gamma$-triggered jets. For the initial hard processes, events were generated with different configurations to isolate specific effects, primarily from the flavor of the parent partons, energy loss, and the selection bias introduced by the $p_{T}$ trigger.
@@
The pgun simulations, in which a single hard parton with specified initial energy and flavor (light quark or gluon) evolves through shower development, revealed a clear flavor dependence in medium modifications to $r_{g}$, $k_{T,g}$, and $m_{g}$, with quark jets exhibiting distinct bump structures that are absent in gluon jets. This contrast arises from the narrow initial structure of quark jets, which allows medium effects to induce larger splittings, whereas the hard, vacuum-like branching of gluon jets remains largely unaffected. The pgun results also highlight the selection bias, which dominates inclusive jet measurements and suppresses jets with broader structures. Lowering the $p^{\mathrm{jet}}_{T}$ cut mitigates this effect, fully eliminating the suppression in $r_{g}$ distributions and partially reducing it in $k_{T,g}$ and $m_{g}$ distributions. These medium modifications were found to originate predominantly at low virtuality ($Q<2$ GeV).
@@
We also presented predictions for $\gamma$-tagged jets by performing simulations with the matter$+$lbt setup within the jetscape framework, incorporating a realistic initial hard process generated by pythia 8. These simulations revealed clear medium-induced broadening in $r_{g}$, $k_{T,g}$, and $m_{g}$. Unlike inclusive jets, which exhibited smooth, monotonic suppression dominated by selection bias, $\gamma$-tagged jets displayed a flat structure at small $r_{g}$ and $k_{T,g}$ and a distinct bump in $m_{g}$, amplified by their quark-jet-dominated nature. Further analysis demonstrated that lowering the $x_{J\gamma}$ cut for $\gamma$-tagged jets effectively reduced selection bias, making the intrinsic modifications more apparent. This study underscores the utility of $\gamma$-tagged jets, and similarly $Z$-tagged jets, as powerful probes for disentangling medium effects on the jet substructure and advancing our understanding of jet-QGP medium interactions in high-energy heavy-ion collisions. When these observables are eventually measured in experiments at high precision, incorporating them into Bayesian analyses with theoretical models could significantly constrain parameters associated with the fundamental mechanisms of jet-QGP medium interactions.
@@ Acknowledgements.
This work was supported in part by the National Science Foundation (NSF) within the framework of the JETSCAPE collaboration, under grant number OAC-2004571 (CSSI:X-SCAPE). It was also supported under PHY-1516590 and PHY-1812431 (R.J.F., M.Ko., C.P. and A.S.); it was supported in part by the US Department of Energy, Office of Science, Office of Nuclear Physics under grant numbers DE-AC02-05CH11231 (X.-N.W. and W.Z.), DE-AC52-07NA27344 (A.A., R.A.S.), DE-SC0013460 (A.K., A.M., C.Sh., I.S., C.Si and R.D.), DE-SC0021969 (C.Sh. and W.Z.), DE-SC0024232 (C.Sh. and H.R.), DE-SC0012704 (B.S.), DE-FG02-92ER40713 (J.H.P. and M.Ke.), DE-FG02-05ER41367 (C.Si, D.S. and S.A.B.), DE-SC0024660 (R.K.E), DE-SC0024347 (J.-F.P. and M.S.). The work was also supported in part by the National Science Foundation of China (NSFC) under grant numbers 11935007, 11861131009 and 11890714 (Y.H. and X.-N.W.), by the Natural Sciences and Engineering Research Council of Canada (C.G., S.J., and G.V.), by the University of Regina President’s Tri-Agency Grant Support Program (G.V.), by the Canada Research Chair program (G.V. and A.K.) reference number CRC-2022-00146, by the Office of the Vice President for Research (OVPR) at Wayne State University (Y.T.), by JSPS KAKENHI Grant No. 22K14041 (Y.T.), and by the São Paulo Research Foundation (FAPESP) under projects 2016/24029-6, 2017/05685-2 and 2018/24720-6 (M.L.). C.Sh., J.-F.P., and R.K.E. acknowledge a DOE Office of Science Early Career Award. I. S. was funded as part of the European Research Council project ERC-2018-ADG-835105 YoctoLHC, and as a part of the Center of Excellence in Quark Matter of the Academy of Finland (project 346325). Calculations for this work used the Wayne State Grid, generously supported by the Office of the Vice President of Research (OVPR) at Wayne State University. The bulk medium simulations were performed using resources provided by the Open Science Grid (OSG) [142, 143], which is supported by the National Science Foundation award #2030508. Data storage was provided in part by the OSIRIS project supported by the National Science Foundation under grant number OAC-1541335.

### v2

@@ V.1 Inclusive jet substructure
In this section, we present the results of groomed observables for inclusive jets. The initial hard processes for the simulations were generated using the pythiagun module with HardQCD:all=on.
@@ V.1.1 Relative transverse momentum of jet splittings
Figure 29: Distributions of the relative transverse momentum of jet splittings, $k_{T,g}$ for charged jets with $60<p^{\mathrm{ch,jet}}_{T}<80$ GeV and $|\eta_{\mathrm{ch,jet}}|<0.7$, using $R=0.2$ in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, and the ratios for different grooming algorithms: soft drop with $z_{\mathrm{cut}}=0.2$ and $\beta=0$ (left), and dynamical grooming with $a=1.0$ (right). The solid lines and circles with statistical error bars show the results from vacuum matter of jetscape and the experimental data from the ALICE Collaboration [18], respectively. The bands indicate the systematic uncertainties of the experimental data.
Figure 30: Ratios of $k_{T,g}$ distributions between Pb-Pb and $p$-$p$ collisions at $\sqrt{s_{NN}}=5.02$ TeV for charged jets with $60<p^{\mathrm{ch,jet}}_{T}<80$ GeV and $|\eta_{\mathrm{ch,jet}}|<0.7$, using $R=0.2$ for different centralities, $0\%$$10\%$ (left) and $30\%$$50\%$ (right), and grooming algorithms, soft drop with $z_{\mathrm{cut}}=0.2$ and $\beta=0$ (upper) and dynamical grooming with $a=1.0$ (lower). The solid lines and squares with statistical error bars show the results from matter+lbt of jetscape and the experimental data from the ALICE Collaboration [18], respectively. The bands indicate the systematic uncertainties of the experimental data.
The distributions of $k_{T,g}$ obtained from soft-drop and dynamical grooming [102], in $p$-$p$ collisions are shown in Fig. 29, and their modifications in Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV are presented in Fig. 30. The results from our jetscape simulations are compared with the experimental results from ALICE [18].
The soft-drop and dynamical grooming results from the JETSCAPE PP19 tune show consistency with experimental data for the $k_{T,g}$ distribution in $p$-$p$ collisions, remaining within the uncertainty range. For Pb-Pb collisions, the modification in the distributions aligns with the inclusive full jet results presented in the main text, demonstrating a monotonically decreasing behavior as a function of $k_{T,g}$, which is primarily governed by selection bias by the $p_{T}$ cut and the loss of $p_{T,2}$. They successfully capture the experimental trends, while dynamical grooming results exhibit slight over-suppression. Notably, the soft-drop results are in excellent agreement with experimental data within the uncertainties.
@@ V.1.2 Groomed jet mass
Figure 31: Distributions of the ratio of the soft-drop groomed jet mass to the jet transverse momentum, $m_{g}/p^{\mathrm{jet}}_{T}$, for full jets with $160<p^{\mathrm{jet}}_{T}<180$ GeV and $|\eta_{\mathrm{jet}}|<1.3$, using $R=0.4$ in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV for different grooming parameters: $z_{\mathrm{cut}}=0.1$ and $\beta=0$ (left), and $z_{\mathrm{cut}}=0.5$ and $\beta=1.5$ (right). The solid lines and circles with statistical error bars show the results from vacuum matter of jetscape and the unfolded experimental data from the CMS Collaboration [123], respectively. The bands indicate the systematic uncertainties of the experimental data.
@@
Figure 32: Distributions of the soft-drop groomed mass, $m_{g}$ for charged jets with $R=0.2$ and $|\eta_{\mathrm{ch,jet}}|<0.7$ in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, and the ratios for different $p^{\mathrm{ch,jet}}_{T}$ range. The soft-drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$. The solid lines and circles with statistical error bars show the results from vacuum matter of jetscape and the experimental data from the ALICE Collaboration [17], respectively. The bands indicate the systematic uncertainties of the experimental data.
@@
Figure 33: Ratios of soft-drop $m_{g}$ distributions between $0\%$$10\%$ Pb-Pb and $p$-$p$ collisions at $\sqrt{s_{NN}}=5.02$ TeV for charged jets with $R=0.2$ and $|\eta_{\mathrm{ch,jet}}|<0.7$ for different $p^{\mathrm{ch,jet}}_{T}$ range. The soft-drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$. The solid lines and squares with statistical error bars show the results from matter+lbt of jetscape and the experimental data from the ALICE Collaboration [17], respectively. The bands indicate the systematic uncertainties of the experimental data.
@@
Here, we present a comparison of our results for soft-drop-groomed jet mass $m_{g}$ with experimental results. The comparison of the unfolded $m_{g}/p_{T}$ distribution for full jets with $R=0.4$ in $p$-$p$ collisions from CMS [123] is shown in Fig. 31, while the comparison of the $m_{g}$ distribution for charged jets with $R=0.2$ in $p$-$p$ collisions from ALICE [17] is shown in Fig. 32. The results for $p$-$p$ collisions are in agreement with the experimental data within the uncertainty range, except for regions with significantly low jet counts, such as the tails of the distributions. It should be noted that, while our calculations are based on Monte Carlo simulations [97], the masses of jet constituents were adjusted to align with experimental methodologies: massless for comparison with CMS data and the charged pion mass for comparison with ALICE data. This adjustment, while necessary to match experimental analysis techniques, has a non-negligible impact on the results, leading to noticeable changes in the distributions.
@@
Figure 33 shows the modifications of the soft-drop-groomed jet mass distributions in $0\%$$10\%$ Pb-Pb collisions, compared with the ALICE data [17]. As with the inclusive full jet results presented in the main text, the behavior is dominated by the $p^{\mathrm{jet}}_{T}$ cut-induced selection bias and the loss of prong mass, leading to suppression at large $m_{g}$. The results capture the experimental trends and agree with the data within the uncertainty range across most $m_{g}$ regions. A slight enhancement observed in some $p^{\mathrm{jet}}_{T}$ ranges for the largest $m_{g}$ bin is primarily attributed to the very low jet counts in the tail of the distribution, particularly in $p$-$p$ collisions.
@@ V.2 Jet splitting radius of $\gamma$-tagged jet
Figure 34: Distributions of the soft-drop-groomed mass, $r_{g}$ for $\gamma$-tagged jets with $R=0.2$ in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, and the ratios for $x_{J\gamma}>0.8$ (left) and $x_{J\gamma}>0.4$ (right). Leading full jets within $|\eta_{\mathrm{jet}}|<2$ and $|\phi_{\mathrm{jet}}-\phi_{\gamma}|>2\pi/3$ are paired with isolated photons satisfying $p^{\gamma}_{T}>100~\mathrm{GeV}$ and $|\eta_{\gamma}|<1.44$ as the trigger condition. The soft-drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$. The solid lines and circles with statistical error bars show the results from vacuum matter of jetscape and the experimental data from the CMS Collaboration [81], respectively. The bands indicate the systematic uncertainties of the experimental data.
Figure 35: Ratios of soft-drop $r_{g}$ distributions between $0\%$$30\%$ Pb-Pb and $p$-$p$ collisions at $\sqrt{s_{NN}}=5.02$ TeV. Full jets of highest $p^{\mathrm{jet}}_{T}$ within $|\eta_{\mathrm{jet}}|<2$ and $|\phi_{\mathrm{jet}}-\phi_{\gamma}|>2\pi/3$ are paired with isolated photons satisfying $p^{\gamma}_{T}>100~\mathrm{GeV}$ and $|\eta_{\gamma}|<1.44$ as the trigger condition. The soft-drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$. The solid lines and squares with statistical error bars show the results from matter+lbt of jetscape and the experimental data from the CMS Collaboration [81], respectively. The bands indicate the systematic uncertainties of the experimental data.
Figures 34 and 35 present comparisons with CMS data [81] for $r_{g}$ distributions in $p$-$p$ collisions and their modifications in $0\%$$30\%$ Pb-Pb collisions, respectively. Results are shown for two lower cuts on $x_{J\gamma}$: $0.8$ and $0.4$. For $p$-$p$ collisions, the distributions obtained from the vacuum matter calculations are narrower than the experimental data for both $x_{J\gamma}$ cut values. This behavior is consistent with the trends observed in other Monte Carlo model calculations compared in Ref. [81].
For the Pb-Pb results, at $x_{J\gamma}>0.8$, suppression at large $r_{g}$ region, attributed to selection bias, is observed, similar to the inclusive jet results presented in the main text and in our previous study [131], as well as the $\gamma$-tagged jet results with large $x_{J\gamma}$ cuts in the main text. While there is a sizable quantitative difference compared with the experimental data, the results qualitatively reproduce the observed trends.
@@ VI Summary
In this work, we investigated medium modifications in jet substructure observables from soft-drop grooming, including the jet splitting momentum fraction ($z_{g}$), splitting radius ($r_{g}$), relative transverse momentum of splittings ($k_{T,g}$), and groomed jet mass ($m_{g}$), in $0\%$$10\%$ Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV. These analyses utilized the matter++lbt multistage jet evolution model within the jetscape framework, using the JETSCAPEv3.5 AA22 tune [88], as employed in previous studies [75, 131]. In particular, this effort should be compared with the closely related earlier companion paper Ref. [121], which studied medium effects on $\gamma$-triggered jets. For the initial hard processes, events were generated with different configurations to isolate specific effects, primarily from the flavor of the parent partons, energy loss, and the selection bias introduced by the $p_{T}$ trigger.
@@
The pgun simulations, in which a single hard parton with specified initial energy and flavor (light quark or gluon) evolves through shower development, reveal a clear flavor dependence in medium modifications to $r_{g}$, $k_{T,g}$, and $m_{g}$. Quark jets exhibit pronounced bump structures absent in gluon jets, with most of the quark-jet bump originating from prongs formed by recoil partons from medium scatterings that satisfy the soft-drop condition and are identified as hard branchings. When recoil effects are artificially disabled, quark jets show only a slight broadening in $r_{g}$ due to other medium effects, such as the transverse diffusion via scatterings and the medium-induced radiation, while the associated branchings remain soft, leading to no visible enhancement in $k_{T,g}$ or $m_{g}$ and resulting in a modification pattern qualitatively identical to gluon jets. The clear flavor dependence arises because quark jets have an intrinsically narrow structure and relatively fewer radiations, making it easier for medium effects to modify the hard prong structures reconstructed by the soft-drop algorithm, whereas the hard, vacuum-like branchings of gluon jets remain largely unaffected. The pgun results also highlight the selection bias, which dominates inclusive jet measurements and suppresses jets with broader structures. Lowering the $p^{\mathrm{jet}}_{T}$ cut mitigates this effect, fully eliminating the suppression in $r_{g}$ distributions and partially reducing it in $k_{T,g}$ and $m_{g}$ distributions. All these medium modifications were found to originate predominantly at low virtuality ($Q<2$ GeV).
@@
We also presented predictions for $\gamma$-tagged jets by performing simulations with the matter++lbt setup within the jetscape framework, incorporating a realistic initial hard process generated by pythia 8. These simulations revealed clear medium-induced modification in $r_{g}$, $k_{T,g}$, and $m_{g}$ of soft-drop-reconstructed hard branchings. Unlike inclusive jets, which exhibited smooth, monotonic suppression dominated by selection bias, $\gamma$-tagged jets displayed a flat structure at small $r_{g}$ and $k_{T,g}$ and a distinct bump in $m_{g}$, amplified by their quark-jet-dominated nature. Further analysis demonstrated that lowering the $x_{J\gamma}$ cut for $\gamma$-tagged jets effectively reduced selection bias, making the intrinsic modifications more apparent.
This study underscores the utility of $\gamma$-tagged jets, and similarly $Z$-tagged jets, as powerful probes for disentangling medium effects on jet substructure and advancing our understanding of jet-QGP interactions in high-energy heavy-ion collisions. Once measured with high precision in experiments, incorporating these observables into Bayesian analyses with theoretical models could provide stringent constraints on the parameters governing the fundamental mechanisms of jet-QGP medium interactions. Furthermore, if a bump-like structure indicating a broadening of the hard substructure is confirmed experimentally, our findings would support the interpretation that the most pronounced medium effects arise primarily from the relatively hard contributions of recoils, making these observables particularly valuable for detailed investigations of recoil dynamics [104, 66, 110].
@@ Acknowledgements.
This work was supported in part by the National Science Foundation (NSF) within the framework of the JETSCAPE collaboration, under grant number OAC-2004571 (CSSI:X-SCAPE). It was also supported under PHY-1516590 and PHY-1812431 (R.J.F., M.Ko., C.P. and A.S.); it was supported in part by the US Department of Energy, Office of Science, Office of Nuclear Physics under Grants No. DE-AC02-05CH11231 (X.-N.W. and W.Z.), No. DE-AC52-07NA27344 (A.A., R.A.S.), No. DE-SC0013460 (A.K., A.M., C.Sh., I.S., C.Si and R.D.), No. DE-SC0021969 (C.Sh. and W.Z.), No. DE-SC0024232 (C.Sh. and H.R.), No. DE-SC0012704 (B.S.), No. DE-FG02-92ER40713 (J.H.P. and M.Ke.), No. DE-FG02-05ER41367 (C.Si, D.S. and S.A.B.), No. DE-SC0024660 (R.K.E), and No. DE-SC0024347 (J.-F.P. and M.S.). The work was also supported in part by the National Science Foundation of China (NSFC) under Grant No. 11935007, No. 11861131009 and No. 11890714 (Y.H. and X.-N.W.), by the Natural Sciences and Engineering Research Council of Canada (C.G., S.J., and G.V.), by the University of Regina President’s Tri-Agency Grant Support Program (G.V.), by the Canada Research Chair program (G.V. and A.K.) reference number CRC-2022-00146, by the Office of the Vice President for Research (OVPR) at Wayne State University (Y.T.), by JSPS KAKENHI Grants No. 22K14041 and No. 25K07303 (Y.T.), and by the São Paulo Research Foundation (FAPESP) under Projects 2016/24029-6, 2017/05685-2 and 2018/24720-6 (M.L.). C.Sh., J.-F.P., and R.K.E. acknowledge a DOE Office of Science Early Career Award. I. S. was funded as part of the European Research Council project ERC-2018-ADG-835105 YoctoLHC, and as a part of the Center of Excellence in Quark Matter of the Academy of Finland (project 346325). Calculations for this work used the Wayne State Grid, generously supported by the Office of the Vice President of Research (OVPR) at Wayne State University. The bulk medium simulations were performed using resources provided by the Open Science Grid (OSG) [Pordes:2007zzb, Sfiligoi:2009cct], which is supported by the National Science Foundation award #2030508. Data storage was provided in part by the OSIRIS project supported by the National Science Foundation under Grant No. OAC-1541335.
DATA AVAILABILITY
All datasets for the current study were generated and analyzed using the open-source software framework jetscapev3.5 with the AA22 tune, available at [2]. The results can be reproduced using the software and methods described in the manuscript.